# Testing the Equivalence Principle on Cosmological Scales Using Peculiar Acceleration Power Spectrum

### v1

@@ Abstract
While the (weak) Equivalence Principle (EP) has been rigorously tested within the solar system, its validity on cosmological scales, particularly in the context of dark matter and dark energy, remains uncertain. In this study, we propose a novel method to test EP on cosmological scales by measuring the peculiar acceleration power spectrum of galaxies using the redshift drift technique. We develop an EP estimator, $E_{\rm ep}$, to evaluate the consistency of the peculiar acceleration power spectrum across different tracers. By calculating the ratio of the peculiar acceleration power spectra of tracers, the ensemble average of $E_{\rm ep}$ is expected to be unity if EP holds on cosmological scales for these tracers. We validate this estimator using N-body simulations, focusing on four redshift bins with $z\leq 1.5$ and scales of $k$ in the range of $0.007$ and $0.32$ $h/\rm Mpc$. By measuring $E_{\rm ep}$ using i) different samples of dark matter particle mock data and ii) low-mass and high-mass halo mock data, we find that the measured $E_{\rm ep}$ values are consistent with unity within the $2\sigma$ level, supporting the validity of $E_{\rm ep}$ on the linear cosmological scales. Taking advantage of advanced observing capabilities, such as next-generation facilities that extend beyond the Square Kilometer Array, the proposed method offers a promising approach for future cosmological tests of EP.
@@ I Introduction
The celebrated general theory of relativity is undoubtedly founded upon the Equivalence Principle (EP), which represents one of its cornerstones. EP states that all non-gravitational phenomena are locally unaffected by gravity, provided that they are performed in a freely falling frame. A specific consequence of this principle is that all objects (including light) fall in accordance with the same laws. Thus far, EP has been subjected to rigorous testing within the solar system (Will, 2014; Touboul et al., 2017). However, its applicability on cosmological scales when considering dark matter and dark energy remains uncertain. In essence, the large-scale structure of the Universe contains valuable information about EP, which can be used to directly evaluate its validity. To date, only indirect methods have been proposed to test EP (Kehagias et al., 2014; Creminelli et al., 2014; Bonvin and Fleury, 2018; Umeh et al., 2021). In this study, we propose a direct test of (weak) EP via the density-weighted peculiar acceleration power spectrum, which can be quantified through the redshift drift technique.
@@
To rigorously test whether EP holds on cosmological scales, it is essential to examine the statistical internal consistency of the peculiar acceleration power spectrum of galaxies. This is a logical consequence of the EP test. As a first step, we propose an EP estimator to evaluate the principle and perform tests on two different mock data sets: dark matter (DM) particles and DM halos. In this study, we explore the potential of radio observations, focusing on an HI galaxy survey beyond the capabilities of the SKA. This survey is expected to cover $10^{8}$ HI galaxies at $z\leq 1.5$ Klöckner et al. (2015); Kang et al. (2025). Moreover, we make a forward-looking assumption that sufficiently long integration times will enable redshift measurements with a precision of $\sim 1~{}{\rm cm/s}$. This precision is significantly higher than the expected capabilities of the SKA redshift drift experiment (Klöckner et al., 2015).
@@ II Methodology
The galaxy acceleration along the line of sight (LOS), analogous to galaxy velocity, can be decomposed into two distinct components: the background cosmic acceleration, $a_{\rm cos}$, and a perturbed component, namely the peculiar acceleration, $a_{\rm pec,\parallel}$. This decomposition can be expressed as
@@
where $H_{0}$ represents the present-day value of the Hubble parameter, $H(z)$. The second term of Eq. 1 accounts for the effect of local gravitational interactions, which cause the object to accelerate relative to the Hubble flow. This effect is determined by the gradient of the gravitational potential ($\phi$):
@@
Note that when applying the redshift drift technique to detect peculiar accelerations, the full expression for the time derivative of galaxy redshift, up to first order in metric perturbations and in $v/c$. In writing Eq. 3, we have made the following simplifications: i) we neglect all perturbative terms related to the local observer (us), as these terms vanish in the power spectrum estimator proposed in this study; ii) we omit two source-related terms: one analogous to the integrated Sachs-Wolfe (ISW) effect and another proportional to the peculiar velocity $v_{\rm pec}$. These terms are subdominant compared to the peculiar acceleration term, as demonstrated in Uzan et al. (2008b). Furthermore, our analysis is conducted in the real space under the plane-parallel limit approximation, though similar conclusions can be extended to the redshift space or angular statistics at linear scales.
@@ II.1 Estimator
$(2\pi)^{3}\delta^{D}(\bm{k}-\bm{k}^{\prime})P_{a_{\parallel}a_{\parallel}}(\bm{k})$ (4)
$=$ $(1+\delta_{g})a_{\parallel}(1+\delta_{g}^{\prime})a_{\parallel}^{\prime}\right>$
$=$ $(1+\delta_{g})a_{{\rm pec},\parallel}(1+\delta_{g}^{\prime})a_{{\rm pec},\parallel}^{\prime}\right>$
$\approx$ $a_{{\rm pec},\parallel}a_{{\rm pec},\parallel}^{\prime}\right>\,.$
where, in the second equality, the constant term $a_{\rm cos}$ contributes only to the $k=0$ mode and is therefore ignored, and $^{\prime}$ means that we ignore the Dirac-delta function in the power spectrum definition. This is because the $k=0$ mode represents the spatially uniform component of the field, which does not carry information about the spatial structure or clustering of galaxies. In the last equality, we restrict our analysis to linear scales and neglect the second-order term $\left\langle\delta_{g}a_{\rm pec,\|}\right\rangle\approx b_{g}\left\langle\delta_{m}a_{\rm pec,\|}\right\rangle$, where $\delta_{g}$ and $\delta_{m}$ are the galaxy and matter density contrasts, respectively, and $b_{g}$ is the galaxy bias. The feasibility of this approximation is validated in Appendix A, where we demonstrate that the second-order term is negligible on linear scales and does not significantly impact our results.
@@
To account for the effects of shot noise, we subtract the shot noise term $\langle a_{{\rm pec},\parallel}^{2}\rangle/n$ from the measured power spectrum, where $n$ represents the number density of a tracer (halos or particles in this study). This subtraction removes discrete sampling effects, yielding the corrected power spectrum:
$(2\pi)^{3}\delta^{D}(\bm{k}-\bm{k}^{\prime})P_{a_{\parallel}a_{\parallel}}(\bm{k})=\left<a_{{\rm pec},\parallel}a_{{\rm pec},\parallel}^{\prime}\right>-\frac{\left<a_{{\rm pec},\parallel}^{2}\right>}{n_{h}}\,,$ (6)
where $a_{{\rm pec},\parallel}^{2}\right>$ is the dispersion of $a_{{\rm pec},\parallel}$ for such tracer.
@@ II.2 Simulation
To validate EP using the estimator $E_{\rm ep}$, we conducted two distinct simulation tests: one using DM particles as tracers of acceleration, and the other using DM halos as tracers. In the first test, we introduced random redshift measurement errors to each DM particle. Given the large number of particles in the simulation, the impact of these errors on the estimator can be well quantified and does not significantly bias the results. In the second test, we assumed an ideal scenario with no redshift errors due to the relatively low number of halos in our simulation. Introducing redshift errors to individual halos would significantly increase uncertainties in the measured acceleration power spectrum, degrading the quality of the EP test.
For both simulation tests, we performed an N-body simulation using the Gadget4 (Springel et al., 2021) code, adopting a $\Lambda$CDM cosmology with parameters: $\sigma_{8}=0.8228$, $n_{s}=0.96$, $\Omega_{m}=0.307115$, $\Omega_{b}=0.049$, and $h=0.68$, compatible with the Planck 2018 measurements within the $2\sigma$ level (Planck Collaboration et al., 2020). The simulation employs a box size of $L_{\rm box}=1024~{}{\rm Mpc}/h$ and a total of $N_{\rm p}=1024^{3}$ particles. Moreover, we employ the Cloud-in-Cell (CIC) method to map the particle/halo acceleration field onto a regular $1024^{3}$ grid.
II.2.1 DM mock generation
To simulate measurement errors in redshift, reflecting a more realistic observational scenario, we introduced redshift errors in each particle. Specifically:
1) For each redshift snapshot, randomly select $10^{8}$ particles from the full set of $1024^{3}$ DM particles in the simulation. This subset is labeled as “group1”. Then, exclude the particles in “group1” from the full set. From the remaining particles, randomly select another $10^{8}$ particles to form “group2”.
@@ II.2.1 DM mock generation
2) Adding random Gaussian redshift noise independently to each particle in both groups. The noise was generated with a mean of zero and a standard deviation of $\sigma_{z_{\parallel}}=f\sigma_{\rm SKA}$, where $\sigma_{\rm SKA}=10~{}{\rm m/s}$ (Klöckner et al., 2015) represents the typical redshift measurement uncertainty expected from SKA observations (an integration time with $\mathcal{O}(1)-\mathcal{O}(10)$ hours per pointing). Here the parameter $f$ is a scaling factor applied to $\sigma_{\rm SKA}$. In this study, we consider two noise levels for enhancing the SKA observation accuracy: Case A with $f=0.001$ and Case B, with $f=\sqrt{2}\times 0.001$, which corresponds to a higher noise level, doubling the noise in the power spectrum.
3) For a given redshift, the acceleration power spectra of the two groups are calculated using Eq. 6, with corrections applied to account for systematic effects. First, we subtract the power spectrum of Poisson shot noise. Additionally, we subtract the mean power spectrum of the random redshift noise, which arises due to the presence of Gaussian-distributed accelerations. This mean noise power spectrum is estimated by computing the power spectrum of $10^{8}$ randomly distributed particles, each assigned Gaussian-distributed accelerations corresponding to either Case A or Case B. The amplitude of this noise power spectrum is approximately $\sim 10^{-16}({\rm Mpc}/h)^{3}{\rm cm^{2}/s^{4}}$.

### v2

@@ Abstract
While the (weak) Equivalence Principle (EP) has been rigorously tested within the solar system, its validity on cosmological scales, particularly in the context of dark matter and dark energy, remains uncertain. In this study, we propose a novel method to test EP on cosmological scales by measuring the peculiar acceleration power spectrum of galaxies using the redshift drift technique. We develop an EP estimator, $E_{\rm ep}$, to evaluate the consistency of the peculiar acceleration power spectrum across different tracers. By calculating the ratio of the peculiar acceleration power spectra of tracers, the ensemble average of $E_{\rm ep}$ is expected to be unity if EP holds on cosmological scales for these tracers. We validate this estimator using N-body simulations, focusing on four redshift bins with $z\leq 1.5$ and scales of $k$ in the range of $0.007$ and $0.2$ $h/\rm Mpc$. By fitting a single parameter $\delta_{\rm ep}$ across redshifts, we find that DM particle mocks without EP violation yield $\delta_{\rm ep}$ consistent with zero under the small redshift measurement uncertainty case, while the large redshift uncertainty case slightly induces biases at low redshifts. In addition, when using DM halo mocks with controlled EP violations, no-violation and mild-violation cases show no significant detection, while moderate and strong violations produce statistically significant $\delta_{\rm ep}$ values and high $\chi^{2}$, especially at low redshifts, confirming the estimator’s sensitivity. Taking advantage of advanced observing capabilities, such as next-generation facilities that extend beyond the Square Kilometer Array, the proposed method offers a promising approach for future cosmological tests of EP.
@@ I Introduction
The celebrated general theory of relativity is undoubtedly founded upon the Equivalence Principle (EP), which represents one of its cornerstones. EP states that all non-gravitational phenomena are locally unaffected by gravity, provided that the experiments are performed in a freely falling frame. A specific consequence of this principle is that all objects (including light) fall in accordance with the same laws. Thus far, EP has been subjected to rigorous testing within the solar system (Will, 2014; Touboul et al., 2017). However, its applicability on cosmological scales when considering dark matter (DM) and dark energy remains uncertain. In essence, the large-scale structure of the Universe contains valuable information about EP, which can be used to directly evaluate its validity. To date, only indirect methods have been proposed to test EP (Kesden and Kamionkowski, 2006a, b; Kehagias et al., 2014; Creminelli et al., 2014; Bonvin and Fleury, 2018; Umeh et al., 2021). In this study, we propose a direct test of (weak) EP via the density-weighted peculiar acceleration power spectrum, which can be quantified through the redshift drift technique.
@@
To rigorously test whether EP holds on cosmological scales, it is essential to examine the statistical internal consistency of the peculiar acceleration power spectrum of galaxies. This is a logical consequence of the EP test. As a first step, we propose in this paper an EP estimator to evaluate the principle and perform tests on two different mock data sets: DM particles and DM halos. In this study, we explore the potential of radio observations, focusing on an HI galaxy survey beyond the capabilities of the SKA. This survey is expected to cover $10^{8}$ HI galaxies at $z\leq 1.5$ Klöckner et al. (2015); Kang et al. (2025). Moreover, we make a forward-looking assumption that sufficiently long integration times will enable redshift measurements with a precision of $\sim 1~{\rm cm/s}$. This precision is significantly higher than the expected capabilities of the SKA redshift drift experiment (Klöckner et al., 2015).
@@ II Methodology
The galaxy acceleration along the line of sight (LOS), analogous to galaxy velocity, can be decomposed into two distinct components: the background cosmic acceleration, $a_{\rm cos}$, and a perturbed component, namely the peculiar acceleration $a_{\rm pec,\parallel}$. This decomposition can be expressed as
@@
where $H_{0}$ represents the present-day value of the Hubble parameter, $H(z)$. This term forms the basis for the redshift drift technique, which is proposed to directly measure the real-time cosmic expansion Sandage (1962); McVittie (1962); Loeb (1998). The second term of Eq. 1 accounts for the effect of local gravitational interactions, which cause the object to accelerate relative to the Hubble flow. This effect is determined by the gradient of the gravitational potential ($\phi$):
@@
Note that when applying the redshift drift technique to detect peculiar accelerations, the full expression for the time derivative of galaxy redshift, up to first order in metric perturbations and in $v_{\rm pec}/c$, was derived in Uzan et al. (2008b). In writing Eq. 3, we have made the following simplifications: i) we neglect all perturbative terms related to the local observer (us), as these terms vanish in the power spectrum estimator proposed in this study; ii) we omit two source-related terms: one analogous to the integrated Sachs-Wolfe (ISW) effect and another proportional to the peculiar velocity $v_{\rm pec}$. These terms are subdominant compared to the peculiar acceleration term, as demonstrated in Uzan et al. (2008b). Furthermore, our analysis is conducted in the 3D redshift space under the plane-parallel limit approximation, though similar conclusions can be extended to 2D angular statistics at linear scales.
@@ II.1 Estimator
$(2\pi)^{3}\delta^{D}(\bm{k}+\bm{k}^{\prime})P_{a_{\parallel}a_{\parallel}}(\bm{k})$ (4)
$=$ $\left<(1+\delta_{g})a_{\parallel}(1+\delta_{g}^{\prime})a_{\parallel}^{\prime}\right$>
$=$ $\left<(1+\delta_{g})a_{{\rm pec},\parallel}(1+\delta_{g}^{\prime})a_{{\rm pec},\parallel}^{\prime}\right$>
$\approx$ $\left<a_{{\rm pec},\parallel}a_{{\rm pec},\parallel}^{\prime}\right$>\,.
where, in the second equality, the constant term $a_{\rm cos}$ contributes only to the $k=0$ mode and is therefore ignored. This is because the $k=0$ mode represents the spatially uniform component of the field, which does not carry information about the spatial structure or clustering of galaxies. In the last equality, we restrict our analysis to linear scales and neglect the second-order term $\left\langle\delta_{g}a_{\rm pec,\|}\right\rangle$ where the galaxy density fluctuation $\delta_{g}$ is much smaller than unity. In Appendix A, we present the theoretical formulation of the linear $P_{a_{\parallel}a_{\parallel}}$ and compare it with simulation measurements to test its validity  [^1] How the violation of this linear approximation will effect the accuracy of the estimator is tested in Sect. III, where we demonstrate that even its breakdown due to non-linearity on small scales does not significantly affect our results..
@@ II.2 Simulation
[…]
To validate the robustness of the estimator $E_{\rm ep}$ on the EP test, we examine three key factors that may affect its measurement: (i) the tracer number density, which influences the shot noise term in Eq. 8; (ii) uncertainties in redshift measurements; and (iii) deviations from the linearity assumption underlying Eq. 5. The first two factors mainly contribute to the statistical uncertainty in the $E_{\rm ep}$ measurement, while the third introduces a potential systematic bias in the EP test.
Ideally, the test should be performed on a simulation combining sufficiently large volume and high mass resolution to both resolve small-mass halos and reproduce the tracer number density expected in upcoming HI surveys like SKA. Due to the limited volume and resolution of our current simulation, we instead employ a two-pronged validation approach: one test utilizes DM particles as tracers to minimize shot noise and isolate the impact of redshift measurement uncertainty on the $E_{\rm ep}$ detection, while the other uses DM halos to assess how shot noise contributes to the statistical uncertainty in the $E_{\rm ep}$ measurement, and how the non-linear evolution of structure formation may introduce systematic bias in the interpretation of $E_{\rm ep}$
For the DM particle test, we treat individual DM particles as acceleration tracers and introduce random redshift measurement errors to mimic observational uncertainties. Thanks to the large number of particles, shot noise is well characterized and does not significantly bias the results. Furthermore, since both tracer populations in this test are drawn from the same underlying DM distribution, the expectation value of $E_{\rm ep}$ is analytically known to be unity and independent of wavenumber $k$. Therefore, this test is free from biases caused by the breakdown of the linear approximation in Eq. 5.
For the halo-based test, we use DM halos as tracers and assume an idealized scenario without redshift errors. This choice is motivated by the fact that redshift uncertainties are inherently significant for the EP test, and including them in a low-number-density sample would introduce substantial noise, severely degrading the precision of the measured acceleration power spectrum. In this test, we focus on the impact of shot noise and the potential breakdown of linearity on small scales. Together, these two complementary setups allow us to isolate and evaluate the effects of tracer density, redshift errors, and nonlinearity on the robustness of the EP estimator.
Beyond quantifying the impact of different sources of uncertainty, we further divide our validation strategy into two physical scenarios based on whether the input tracer samples contain intrinsic EP violation. In the first scenario, the samples are constructed to obey EP, allowing us to evaluate how tracer number density, redshift uncertainties, and nonlinearity affect the signal-to-noise ratio of the $E_{\rm ep}$ measurement. In the second scenario, we introduce artificial EP violations into the halo catalogs by modifying the acceleration field in a controlled manner. This enables a direct assessment of the sensitivity and robustness of the $E_{\rm ep}$ estimator in detecting potential deviations from EP.
II.2.1 DM particle mock generation
We first test EP violation using the DM particle mock. To calculate the estimator $E_{\rm ep}$, which is defined as the ratio of power spectra from two distinct datasets, we proceed as follows. For each redshift snapshot, we randomly select $10^{8}$ particles from the full set of $1024^{3}$ DM particles in the simulation to form the first dataset. These selected particles are then excluded from the full set, and another $10^{8}$ particles are randomly chosen from the remaining particles to construct the second dataset.
@@ II.2.1 DM particle mock generation
To mimic realistic observational uncertainties, we simulate redshift measurement errors by adding independent Gaussian noise to the redshift of each particle in both datasets. The noise was generated with a mean of zero and a standard deviation of $\sigma_{z_{\parallel}}=f\sigma_{\rm SKA}/t_{\rm obs}$, where $\sigma_{\rm SKA}=10~{\rm m/s}$ (Klöckner et al., 2015) represents the typical redshift measurement uncertainty expected from SKA observations (an integration time with $\mathcal{O}(1)-\mathcal{O}(10)$ hours per pointing), and $t_{\rm obs}=12$ years. Here the parameter $f$ is a scaling factor applied to $\sigma_{\rm SKA}$. In this study, we consider two noise levels to assess their impact on SKA observation accuracy: a low-noise case with $f=0.001$, and a high-noise case with $f=0.002$. The latter corresponds to a doubled noise amplitude, resulting in four times higher noise power in the power spectrum.
II.2.2 DM Halo Mock generation
$z$ $N_{h}$ (small/large) $M_{h}[10^{12}~M_{\odot}/h$] (small/large)
1.5 $360821/1082464$ [2.733.41]/[3.41847.050][2.73-3.41]/[3.41-847.050]
1.0 $466526/1399581$ [2.733.49]/[3.491387.33][2.73-3.49]/[3.49-1387.33]
0.5 $536364/1609094$ [2.733.58]/[3.582421.35][2.73-3.58]/[3.58-2421.35]
0.0 $550123/1650372$ [2.733.66]/[3.665085.91][2.73-3.66]/[3.66-5085.91]
Table 1: Number of halos in the “small-mass” and “large-mass” groups, along with their corresponding mass ranges, are summarized at different redshifts.
In addition to testing EP using DM particle simulation data, we also conduct the second test by comparing the acceleration power spectra of DM halos with different masses. In the simulation, DM halos are identified using the group finder ROCKSTAR (Behroozi et al., 2013). We selected all halos containing at least ten DM particles, corresponding to $M_{h}\geq 2.73\times 10^{12}~M_{\odot}/h$. These halos are subsequently divided into two mass bins, referred to as “small-mass” and “large-mass”, with the former containing halos of smaller mass and the latter containing halos of larger mass (Tab. 1). For each fixed redshift, the ratio of the number of small-mass halos to large-mass halos is maintained at approximately $1/3$. This ratio ensures that the shot noise contributions from the small and large mass bins are nearly equal, as confirmed by our measurements.
To assess the sensitivity of our estimator to violations of EP, we introduce artificial EP-breaking effects into the halo catalogs as follows. For each halo of mass $M$, we rescale its line-of-sight (LOS) acceleration by a mass-dependent factor $\lambda$:
$a_{\parallel,\mathrm{art}}=\lambda a_{\parallel},\quad\text{with}\quad\lambda=1+\alpha\left(\frac{M}{M_{\ast}}\right)^{\beta}\,.$ (6)
This model is not derived from any specific underlying theory, but instead serves as a controlled, parameterized prescription to introduce EP violations as a function of halo mass. It is a purely phenomenological framework, designed to systematically probe the sensitivity of the estimator to such violations. When $\alpha=0$, the EP is exactly preserved. Nonzero values of $\alpha$ introduce controlled deviations from EP. In this study, the characteristic mass scale is fixed at $M_{*}=10^{12}M_{\odot}$, as it represents the characteristic mass of typical galaxies.
We fix $\alpha=0.1$ and vary $\beta$ to control the mass dependence of the EP violation. Within this phenomenological framework, we consider representative values of $\beta=0.1$, $0.3$, and $0.5$ to assess how different scalings of the EP-breaking term influence the response of our estimator. Under these parameter choices, the relative deviations of $a_{||}$ from $\lambda$ in the small-mass group reach approximately 10-15% for $\beta=0.1$ and $0.3$, and up to 18% for $\beta=0.5$. In contrast, for the large-mass group, the deviations are more significant–about 20% for $\beta=0.1$, 100% for $\beta=0.3$, and a factor of 3.9–8.1 for $\beta=0.5$.
II.3 Measurements of $P_{a_{\parallel}a_{\parallel}}$ and its covariance
The method by which we measure the power spectrum of the LOS acceleration, $P_{a_{\parallel}a_{\parallel}}$, from both observational and simulation data is as follows. In observations, we employ the redshift drift technique to infer the LOS acceleration $a_{||}$ via the relation $\dot{z}=a_{||}/c$ for each galaxy. In simulations, we directly extract the acceleration of DM particles or halos from the simulation snapshots. In particular, the acceleration of a halo is that of the particle nearest to the halo center.
The computation of $P_{a_{\parallel}a_{\parallel}}$ follows an approach analogous to the standard method used for estimating the power spectrum of the mass-weighted velocity field (i.e., the momentum field). Specifically, we proceed as follows:
  1. 1.
    We define a cubical volume enclosing the full survey or simulation region, as described in Sect. II.2. For simulations, this naturally corresponds to the simulation box itself. A regular grid is then constructed within this volume.
  2. 2.
    Using the Cloud-in-Cell (CIC) method, we interpolate the mass-weighted acceleration of tracers (galaxies, halos, or DM particles) onto the grid to construct a continuous acceleration field. For simulation data, prior to this assignment, we convert tracer positions from real space to redshift space along a chosen LOS direction, i.e., the $x$-axis, as:
    $x_{s}=x_{r}+\frac{v_{x}}{aH(z)}\,.$ (7)
    Here, $x_{s}$ and $x_{r}$ denote the redshift-space and real-space positions along LOS, respectively; $v_{x}$ is the LOS component of the proper peculiar velocity; $a$ is the scale factor; and $H(z)$ is the Hubble parameter at redshift $z$.
  3. 3.
    We perform a Fast Fourier Transform (FFT) on the gridded, mass-weighted LOS acceleration field to obtain its Fourier components. The power spectrum $P_{a_{\parallel}a_{\parallel}}(k)$ is then computed by averaging the squared amplitude of the Fourier modes over spherical shells in $k$-space, i.e., $P_{a_{\parallel}a_{\parallel}}(k)=\langle|a_{{\rm pec},\parallel}(\mathbf{k})|^{2}\rangle_{|\mathbf{k}|=k}$. This yields the isotropic (angle-averaged) component of the power spectrum. Accordingly, in this study we focus exclusively on the monopole of $P_{a_{\parallel}a_{\parallel}}$, leaving any anisotropic features for future investigation.
  4. 4.
    After measuring the power spectrum $P_{a_{\parallel}a_{\parallel}}$, we apply corrections for systematic effects arising from shot noise and redshift measurement uncertainties. First, we subtract the Poisson shot noise contribution, estimated by $\langle a^{2}_{\rm pec,\parallel}/n_{h}\rangle$, such as
    $(2\pi)^{3}\delta^{D}(\bm{k}+\bm{k}^{\prime})P_{a_{\parallel}a_{\parallel}}(\bm{k})$ (8)
    $=$ $\left<a_{{\rm pec},\parallel}a_{{\rm pec},\parallel}^{\prime}\right>-\frac{\left<a_{{\rm pec},\parallel}^{2}\right>}{n_{h}}\,$,
    where $n_{h}$ represents the number density of the tracer (halos or particles in this study), and $\left<a_{{\rm pec},\parallel}^{2}\right$> is the dispersion of $a_{{\rm pec},\parallel}$ for such tracer. For the DM mocks, we further correct for the additional noise induced by redshift measurement errors. This redshift uncertainty introduces Gaussian-distributed noise in the acceleration measurements, effectively contributing an extra shot-noise-like component to the power spectrum. We estimate the impact by assigning Gaussian acceleration noise (low or high) to $10^{8}$ randomly distributed particles, yielding a noise power spectrum with amplitude $\sim 10^{-16}\,({\rm Mpc}/h)^{3}\,{\rm cm}^{2}/{\rm s}^{4}$.